VIBRONIC ANALYSIS OF THE Ã^{2}E′′ STATE OF NO_{3} RADICAL

\underline{TERRANCE J. CODD}, MOURAD ROUDJANE, MING-WEI CHEN^{a}, and TERRY A. MILLER, Laser Spectroscopy Facility, The Ohio State University, Columbus, Ohio 43210.

The nitrate radical is a key reactant in atmospheric chemistry leading to the formation of acid rain and is the primary oxidant in the night sky. The Ã^{2}E′′ state of NO_{3} is doubly degenerate and is therefore subject to Jahn-Teller (JT) coupling through the degenerate in-plane stretch and bend modes (ν_{3} and ν_{4} respectively). We have taken a moderate resolution CRDS spectrum of the Ã^{2}E′′-X̃^{2}A′_{2} transition of the NO_{3} radical under jet-cooled conditions. We resolve ∼ 20 vibronic transitions and are able to assign many using an independent anharmonic oscillator model as was presented previously.^{b} In order to gain a deeper understanding of the nature of the JT effect in this electronic state we have performed a vibronic analysis including linear and quadratic JT coupling terms for ν_{3} and ν_{4} and possible bilinear coupling between the totally symmetric stretch, ν_{1}, and ν_{4}. We conclude that the JT coupling in ν_{4} is quite weak. Satisfactory spectral fits can be obtained assuming weak JT coupling for ν_{3} also, though there is some evidence of strong JT coupling for ν_{3} and the strengths and weaknesses of each case are discussed.

aPresent address: University of Illinois at Urbana-Champaign, Urbana, IL 61801

bCodd, T. et al. 67^{th} Int. Symp. Molec. Spec. (2012)